rxcklesskaisher:

If the area of the rectangle to be drawn is 42 square units, where should points B and D be located if they lie to the right of points A and C?

rxcklesskaisher:

1 attachment
JamesTDG:

Try to solve for a=(b*h)/2 to find the coordinates of the tip

rxcklesskaisher:

A?

JamesTDG:

a is for area, b is for base, h is for height

rxcklesskaisher:

B(4, 4) and D(4, −3) B(−3, 4) and D(4, 4) B(2, 3) and D(2, −4) B(2, 4) and D(2, −3)

rxcklesskaisher:

choices

JamesTDG:

odd

surjithayer:

diffrence between A and C is 7 area of rectangle=length *width. Here width=7 squares length=area/width=42/7=6 squares. so B and D should be 2 squares right of Y-axis . so B and D are (2,4) and (2,-3)

rxcklesskaisher:

so?

rxcklesskaisher:

@surjithayer